OpenStudy (anonymous):

y-2x=3 3x-2y=5 Solving System of Equation Using Substitution

OpenStudy (anonymous):

Need to get the 'x' and 'y' on the same side

OpenStudy (anonymous):

Substitution is perfect here because you already have a variable isolated on one side. y =8-x, right? So, you can just substitute (8-x) for y in the second equation. 7 = 2-y 7 = 2-(8-x) Then you go from there until you find x. Once you find x, you put its value in the first equation to find y, and the solution of the system is (x,y) @T0mmy that isn't substitution!

OpenStudy (anonymous):

Thanxs

OpenStudy (anonymous):

Yep my mistake sorry.